OpenStudy (anonymous):

Find g[h(x)] and h[g(x)] Given: g(x)=4x h(x)=2x-1

myininaya (myininaya):

\[g[h(x)]=g[2x-1]=4(2x-1)\] \[h[g(x)]=h[4x]=2(4x)-1\]

OpenStudy (anonymous):

okay i see what you did there..i understand now! thanks myin! oh and happy belated birthday!!

myininaya (myininaya):

it was yesterday but thanks :)